A rectangular plot of ground measuring 12 meters by 20 meters is surrounded by a sidewalk of uniform width. The area of the sidewalk is 68 square meters. Find the width of the walk.
Answer by Paul(988)

The width is 12
Let the width be x
x-12
and
x+20

(x-12)(x+20)=68
x^2+20x-12x-240=68
x^2+8x-308=0,x

Factor:
(x+22) and (x-14)

hence, the width is 14m.
Paul.